OpenStudy (anonymous):

One number is 2 more than 5 times another. Their product is 72. Find the numbers.

OpenStudy (anonymous):

let the number be a and b a=2+5b a*b=72 solve these equations

OpenStudy (anonymous):

x+2=5y x*y=72 x=5y-2 plug into the other (5y-2)*y=72 5y^2-2y-72=0 (5y+18)(y-4)=0 5y=-18 y=-18/5 and y-4=0 y=4 (x+2)=5(-18/5) and (x+2)=5(4) x=-20 and x=18